# A Survey on Occupancy Perception for Autonomous Driving: The Information Fusion Perspective

### v1

@@ Abstract
3D occupancy perception technology aims to observe and understand dense 3D environments for autonomous vehicles. Owing to its comprehensive perception capability, this technology is emerging as a trend in autonomous driving perception systems, and is attracting significant attention from both industry and academia. Similar to traditional bird’s-eye view (BEV) perception, 3D occupancy perception has the nature of multi-source input and the necessity for information fusion. However, the difference is that it captures vertical structures that are ignored by 2D BEV. In this survey, we review the most recent works on 3D occupancy perception, and provide in-depth analyses of methodologies with various input modalities. Specifically, we summarize general network pipelines, highlight information fusion techniques, and discuss effective network training. We evaluate and analyze the occupancy perception performance of the state-of-the-art on the most popular datasets. Furthermore, challenges and future research directions are discussed. We hope this report will inspire the community and encourage more research work on 3D occupancy perception. A comprehensive list of studies in this survey is available in an active repository that continuously collects the latest work: https://github.com/HuaiyuanXu/3D-Occupancy-Perception.
@@
Autonomous Driving, Information Fusion, Occupancy Perception.
@@ 1.1 Occupancy Perception in Autonomous Driving
Autonomous driving can improve urban traffic efficiency and reduce energy consumption. For reliable and safe autonomous driving, a crucial capability is to accurately and comprehensively understand the surrounding environment, that is, to perceive the observed world. At present, bird’s-eye view (BEV) perception is the mainstream perception pattern [1, 2], with the advantages of absolute scale and no occlusion for describing environments. BEV perception provides a unified representation space for multi-source information fusion (e.g., information from diverse viewpoints, modalities, sensors, and time series) and numerous downstream applications (e.g., explainable decision making and motion planning). However, BEV perception does not monitor height information, thereby cannot provide a complete representation for the 3D scene.
To address this, occupancy perception was proposed for autonomous driving to capture the dense 3D structure of the real world. This burgeoning perception technology aims to infer the occupied state of each voxel for the voxelized world, characterized by a strong generalization capability to open-set objects, irregular-shaped vehicles, and special road structures [3, 4]. Compared with 2D views such as perspective view and bird’s-eye view, occupancy perception has a nature of 3D attributes, making it more suitable for 3D downstream tasks, such as 3D detection [5, 6] and segmentation [4].
Figure 1: Autonomous driving vehicle system. The sensing data from cameras, LiDAR, and radar enable the vehicle to intelligently perceive its surroundings. Subsequently, the intelligent decision module generates control and planning of driving behavior. Occupancy perception surpasses the perception based on the perspective view, bird’s-eye view, or point-cloud space, in terms of 3D understanding, density, and no occlusion.
In academia and industry, occupancy perception for holistic 3D scene understanding poses a meaningful impact. On the academic consideration, it is challenging to estimate dense voxel-wise occupancy of the real 3D world from complex input formats, encompassing multiple sensors, modalities, and temporal sequences. Moreover, it is valuable to further reason about semantic categories [7], textual descriptions [8], and motion states [9] for occupied voxels, which paves the way toward a more comprehensive understanding of the environment. From the industrial perspective, the deployment of a LiDAR Kit on each autonomous vehicle is expensive. With cameras as a cheap alternative to LiDAR, vision-centric occupancy perception is indeed a cost-effective solution that reduces the manufacturing cost for vehicle equipment manufacturers.
@@ 1.2 Motivation to Information Fusion Research
The gist of occupancy perception lies in comprehending complete and dense 3D scenes, including understanding occluded areas. However, the observation from a single sensor, such as a 2D image or a point cloud, only captures parts of the scene. Figure 1 intuitively illustrates that an image or a point cloud cannot provide a 3D panorama or a dense environmental scan, resulting in inferior scene perception. To this end, studying information fusion from multiple sensors [10, 11, 12] and multiple frames [4, 7] will facilitate the comprehensive occupancy perception. This is because, on the one hand, information fusion expands the range of spatial perception, and on the other hand, it densifies scene observation. Besides, for occluded regions, integrating multi-frame observations is beneficial, as the same scene is observed by a host of viewpoints, offering sufficient scene features for occlusion inference.
Furthermore, in dynamic outdoor scenarios, where autonomous vehicles must navigate complex environments under varying lighting and weather conditions, the need for stable occupancy perception is paramount. The robustness of perception is crucial for ensuring driving safety and efficiency. At this point, research on multi-modal fusion will promote robust occupancy perception, by combining the strengths of different modalities of data [10, 11, 13, 14]. For example, LiDAR and radar data are insensitive to illumination changes and can sense the precise depth of the scene. This capability is particularly important during nighttime driving or in scenarios where shadows and glare could obscure critical information. Camera data excel in capturing detailed visual texture, being adept at identifying long-distance and color-based environmental elements, such as road signs, traffic lights, and lane markings. The fusion of data from these diverse modalities will present a holistic view of the environment meanwhile against adverse environmental changes.
@@ 1.3 Contributions
On related perception topics, 3D semantic segmentation [15, 16] and 3D object detection [17, 18, 19] have been extensively reviewed. However, these tasks do not facilitate a dense understanding of the environment. BEV perception, which addresses this issue, has also been thoroughly reviewed [2, 1]. Our survey focuses on 3D occupancy perception, which captures the environmental height information overlooked by BEV perception. Roldao et al. [20] conducted a literature review on 3D scene completion for indoor and outdoor scenes, which is closely related to our focus. Unlike their work, our survey is specifically tailored to autonomous driving scenarios. Moreover, given the multi-source nature of 3D occupancy perception, we provide an in-depth analysis of information fusion techniques for this field. The primary contributions of this survey are three-fold:
@@
    We systematically review the latest research on 3D occupancy perception in the field of autonomous driving, covering the overall research background, comprehensive analyses of its importance, and in-depth discussions on related techniques.
@@
    We provide the taxonomy of 3D occupancy perception, and elaborate on core methodological issues, including network pipelines, multi-source information fusion, and effective network training.
@@ 2.1 A Brief History of Occupancy Perception
Occupancy perception is derived from Occupancy Grid Mapping (OGM) [21], which is a classic topic in mobile robot navigation, and aims to generate a grid map from noisy and uncertain measurements. Each grid in this map is assigned a value that scores the probability of the grid space being occupied by obstacles. Semantic occupancy perception originates from SUNCG [22], which predicts the occupied status and semantics of all voxels in an indoor scene from a single image. However, studying occupancy perception in outdoor scenes is imperative for autonomous driving, as opposed to indoor scenes. MonoScene [23] is a pioneering work of outdoor scene occupancy perception using only a monocular camera. Contemporary with MonoScene, Tesla announced its brand-new camera-only occupancy network at the CVPR 2022 workshop on Autonomous Driving [24]. This new network comprehensively understands the 3D environment surrounding a vehicle according to surround-view RGB images. Subsequently, occupancy perception has attracted extensive attention, catalyzing a surge in research on occupancy perception for autonomous driving in recent years.
@@ 2.3.1 Bird’s-Eye-View Perception
3D reconstruction is a traditional but important topic in the computer vision and robotics communities. The objective of 3D reconstruction from images is to construct 3D of an object or scene based on 2D images captured from one or more viewpoints. Early methods exploited shape-from-shading [60] or structure-from-motion [61]. Afterwards, the neural radiation field (NeRF) [62] introduced a novel paradigm for 3D reconstruction, which learned the density and color fields of 3D scenes, producing results with unprecedented detail and fidelity. However, such performance necessitates substantial training time and resources for rendering [63, 64, 65], especially for high-resolution output. Recently, 3D Gaussian splatting (3D GS) [66] has addressed this issue by redefining a paradigm-shifting approach to scene representation and rendering. Specifically, it represents scene representation with millions of 3D Gaussian functions in an explicit way, achieving faster and more efficient rendering [67]. 3D reconstruction emphasizes the geometric quality and visual appearance of the scene. In comparison, voxel-wise occupancy perception has lower resolution and visual appearance requirements, focusing instead on the occupancy distribution and semantic understanding of the scene.
@@ 3 Methodologies
Figure 3: Architectures for occupancy perception. For compactness, we do not display all the details but focus on the overall pipelines of LiDAR-centric, vision-centric, and multi-modal occupancy perception approaches.
@@ 3.1.1 General Pipeline
Fig. 3(a) illustrates the general pipeline of LiDAR-centric occupancy perception. The input point cloud first undergoes feature extraction and voxelization, followed by representation enhancement via an encoder-decoder module. Ultimately, the complete and dense occupancy of the scene is inferred. Specifically, given a point cloud $P\in\mathbb{R}^{N\times 3}$ and obtaining the features $F\in\mathbb{R}^{N\times C}$ of each point, a series of initial voxels are produced. These voxels are distributed in a 3D volume [25, 74, 27], a 2D BEV plane [68, 27], or three 2D tri-perspective view planes [72]. This operation constructs the 3D feature volume or 2D feature map, denoted as $V_{\text{init}-3D}\in\mathbb{R}^{X\times Y\times Z\times D}$ and $V_{\text{init}-2D}\in\mathbb{R}^{X\times Y\times D}$ respectively. $N$ represents the number of points; $\left(X,Y,Z\right)$ are the length, width, and height; $C$ and $D$ mean the feature dimensions of points and voxels, respectively. In addition to voxelizing in regular Euclidean space, PointOcc [72] builds tri-perspective 2D feature maps in a cylindrical coordinate system. The cylindrical coordinate system aligns more closely with the spatial distribution of points in the LiDAR point cloud, where points closer to the LiDAR sensor are denser than those at farther distances. Therefore, it is reasonable to use smaller-sized cylindrical voxels for fine-grained modeling in nearby areas. The feature extraction and voxelization of point clouds can be formulated as:
$V_{\text{init}-2D\text{/}3D}=\Psi_{V}\left(\Phi_{P}\left(P\right)\right),$ (3)
where $\Phi_{P}$ is the feature extractor such as extracting neural features [93] or directly counting the geometric features of points within the voxel (e.g., mean, minimum, and maximum heights) [27]; $\Psi_{V}$ stands for pillar or cubic voxelization. Encoder and decoder can be various modules to enhance features. The final 3D occupancy inference involves applying convolution [25, 27, 71] or MLP [68, 72, 74] on the enhanced features to infer the occupied status $\left\{1,0\right\}$ of each voxel, and even estimate its semantic category:
@@ 3.2.1 General Pipeline
Inspired by Tesla’s technology of the perception system for their autonomous vehicles [24], vision-centric occupancy perception has garnered significant attention both in industry and academia. Compared to LiDAR-centric methods, vision-centric occupancy perception, which solely relies on camera sensors, represents a current trend. There are three main reasons: (i) Cameras are cost-effective for large-scale deployment on vehicles. (ii) RGB images capture rich environmental textures, aiding in the understanding of scenes and objects such as traffic signs and lane lines. (iii) The burgeoning advancement of deep learning technologies enables a possibility to achieve 3D occupancy perception from 2D vision. Vision-centric occupancy perception can be divided into monocular solutions [97, 51, 23, 48, 49, 30, 52, 82, 78] and multi-camera solutions [50, 98, 28, 35, 58, 73, 95, 29, 75, 7]. Multi-camera perception, which covers a broader field of view, follows a general pipeline as shown in Fig. 3(b). It begins by extracting front-view feature maps from multi-camera images, followed by a 2D-to-3D transformation, spatial information fusion, and optional temporal information fusion, culminating with an occupancy head that infers the environmental 3D occupancy.
@@ 3.2.3 Information Fusion in Vision-Centric Occupancy
Sec. 3.2.2 discusses the transformation from a single front-view feature map to a 3D feature volume. In a multi-camera setting, each camera’s front-view feature map describes a part of the scene. To comprehensively understand the scene, it is necessary to spatially fuse the information from multiple feature maps. Additionally, objects in the scene might be occluded or in motion. Temporally fusing feature maps of multiple frames can help reason about the occluded areas and recognize the motion status of objects.
@@
Once the alignment is completed, the historical and current features are fed to the feature fusion module to enhance the representation, especially to strengthen the reasoning ability for occlusion and the recognition ability of moving objects. There are three main streamlines to feature fusion, namely convolution, cross attention, and adaptive mixing. PanoOcc [4] concatenates the previous features with the current ones, then fuses them using a set of 3D residual convolution blocks. Many occupancy perception methods [30, 53, 20, 78] utilize cross attention for fusion. The process is similar to multi-camera cross attention (refer to Eq. 8), but the difference is that 3D-space voxels are projected to 2D multi-frame feature maps instead of multi-camera feature maps. Moreover, FullySparse [96] employs adaptive mixing [105] for temporal information fusion. For the query feature of the current frame, FullySparse samples $S_{n}$ features from historical frames, and aggregates them through adaptive mixing. Specifically, the sampled features are multiplied by the channel mixing matrix $W_{C}$ and the point mixing matrix $W_{S_{n}}$, respectively. These mixing matrices are dynamically generated from the query feature $F_{q}$ of the current frame:
@@
The features resulting from spatial and temporal information fusion are processed by various types of heads to determine 3D occupancy. These include convolutional heads, mask decoder heads, linear projection heads, and linear projection with threshold heads. Convolution-based heads [9, 58, 35, 98, 69, 23] consist of multiple 3D convolutional layers. Mask decoder-based heads [79, 52, 96, 84], inspired by MaskFormer [106] and Mask2Former [107], formalize 3D semantic occupancy prediction into the estimation of a set of binary 3D masks, each associated with a corresponding semantic category. Specifically, they compute per-voxel embeddings and assess per-query embeddings along with their related semantics. The final occupancy predictions are obtained by calculating the dot product of these two embeddings. Linear projection-based heads [4, 29, 33, 48, 78, 30, 83] leverage lightweight MLPs on the dimension of feature channels to produce occupied status and semantics. Furthermore, for the occupancy methods[77, 75, 81, 89, 85] based on NeRF [62], their occupancy heads use two separate MLPs ($\text{MLP}_{\sigma}$, $\text{MLP}_{s}$) to estimate the density volume $V_{\sigma}$ and the semantic volume $V_{S}$. Then the occupied voxels are selected based on a given confidence threshold $\tau$, and their semantic categories are determined based on $V_{S}$:

### v2

@@ Abstract
3D occupancy perception technology aims to observe and understand dense 3D environments for autonomous vehicles. Owing to its comprehensive perception capability, this technology is emerging as a trend in autonomous driving perception systems, and is attracting significant attention from both industry and academia. Similar to traditional bird’s-eye view (BEV) perception, 3D occupancy perception has the nature of multi-source input and the necessity for information fusion. However, the difference is that it captures vertical structures that are ignored by 2D BEV. In this survey, we review the most recent works on 3D occupancy perception, and provide in-depth analyses of methodologies with various input modalities. Specifically, we summarize general network pipelines, highlight information fusion techniques, and discuss effective network training. We evaluate and analyze the occupancy perception performance of the state-of-the-art on the most popular datasets. Furthermore, challenges and future research directions are discussed. We hope this paper will inspire the community and encourage more research work on 3D occupancy perception. A comprehensive list of studies in this survey is publicly available in an active repository that continuously collects the latest work: https://github.com/HuaiyuanXu/3D-Occupancy-Perception.
@@
Autonomous Driving, Information Fusion, Occupancy Perception, Multi-Modal Data.
@@ 1.1 Occupancy Perception in Autonomous Driving
Autonomous driving can improve urban traffic efficiency and reduce energy consumption. For reliable and safe autonomous driving, a crucial capability is to understand the surrounding environment, that is, to perceive the observed world. At present, bird’s-eye view (BEV) perception is the mainstream perception pattern [1, 2], with the advantages of absolute scale and no occlusion for describing environments. BEV perception provides a unified representation space for multi-source information fusion (e.g., information from diverse viewpoints, modalities, sensors, and time series) and numerous downstream applications (e.g., explainable decision making and motion planning). However, BEV perception does not monitor height information, thereby cannot provide a complete representation for the 3D scene.
To address this, occupancy perception was proposed for autonomous driving to capture the dense 3D structure of the real world. This burgeoning perception technology aims to infer the occupied state of each voxel for the voxelized world, characterized by a strong generalization capability to open-set objects, irregular-shaped vehicles, and special road structures [3, 4]. Compared with 2D views such as perspective view and bird’s-eye view, occupancy perception has a nature of 3D attributes, making it more suitable for 3D downstream tasks, such as 3D detection [5, 6], segmentation [4], and tracking [7].
Figure 1: Autonomous driving vehicle system. The sensing data from cameras, LiDAR, and radar enable the vehicle to intelligently perceive its surroundings. Subsequently, the intelligent decision module generates control and planning of driving behavior. Occupancy perception surpasses other perception methods based on perspective view, bird’s-eye view, or point clouds, in terms of 3D understanding and density.
In academia and industry, occupancy perception for holistic 3D scene understanding poses a meaningful impact. On the academic consideration, it is challenging to estimate dense 3D occupancy of the real world from complex input formats, encompassing multiple sensors, modalities, and temporal sequences. Moreover, it is valuable to further reason about semantic categories [8], textual descriptions [9], and motion states [10] for occupied voxels, which paves the way toward a more comprehensive understanding of the environment. From the industrial perspective, the deployment of a LiDAR Kit on each autonomous vehicle is expensive. With cameras as a cheap alternative to LiDAR, vision-centric occupancy perception is indeed a cost-effective solution that reduces the manufacturing cost for vehicle equipment manufacturers.
@@ 1.2 Motivation to Information Fusion Research
The gist of occupancy perception lies in comprehending complete and dense 3D scenes, including understanding occluded areas. However, the observation from a single sensor only captures parts of the scene. For instance, Fig. 1 intuitively illustrates that an image or a point cloud cannot provide a 3D panorama or a dense environmental scan. To this end, studying information fusion from multiple sensors [11, 12, 13] and multiple frames [4, 8] will facilitate a more comprehensive perception. This is because, on the one hand, information fusion expands the spatial range of perception, and on the other hand, it densifies scene observation. Besides, for occluded regions, integrating multi-frame observations is beneficial, as the same scene is observed by a host of viewpoints, which offer sufficient scene features for occlusion inference.
Furthermore, in complex outdoor scenarios with varying lighting and weather conditions, the need for stable occupancy perception is paramount. This stability is crucial for ensuring driving safety. At this point, research on multi-modal fusion will promote robust occupancy perception, by combining the strengths of different modalities of data [11, 12, 14, 15]. For example, LiDAR and radar data are insensitive to illumination changes and can sense the precise depth of the scene. This capability is particularly important during nighttime driving or in scenarios where the shadow/glare obscures critical information. Camera data excel in capturing detailed visual texture, being adept at identifying color-based environmental elements (e.g., road signs and traffic lights) and long-distance objects. Therefore, the fusion of data from LiDAR, radar, and camera will present a holistic understanding of the environment meanwhile against adverse environmental changes.
Figure 2: Chronological overview of 3D occupancy perception. It can be observed that: (1) research on occupancy has undergone explosive growth since 2023; (2) the predominant trend focuses on vision-centric occupancy, supplemented by LiDAR-centric and multi-modal methods.
@@ 1.3 Contributions
Among perception-related topics, 3D semantic segmentation [16, 17] and 3D object detection [18, 19, 20, 21] have been extensively reviewed. However, these tasks do not facilitate a dense understanding of the environment. BEV perception, which addresses this issue, has also been thoroughly reviewed [2, 1]. Our survey focuses on 3D occupancy perception, which captures the environmental height information overlooked by BEV perception. Roldao et al. [22] conducted a literature review on 3D scene completion for indoor and outdoor scenes, which is close to our focus. Unlike their work, our survey is specifically tailored to autonomous driving scenarios. Moreover, given the multi-source nature of 3D occupancy perception, we provide an in-depth analysis of information fusion techniques for this field. The primary contributions of this survey are three-fold:
@@
    We systematically review the latest research on 3D occupancy perception in the field of autonomous driving, covering motivation analysis, the overall research background, and an in-depth discussion on methodology, evaluation, and challenges.
@@
    We provide a taxonomy of 3D occupancy perception, and elaborate on core methodological issues, including network pipelines, multi-source information fusion, and effective network training.
@@ 2.1 A Brief History of Occupancy Perception
Occupancy perception is derived from Occupancy Grid Mapping (OGM) [23], which is a classic topic in mobile robot navigation, and aims to generate a grid map from noisy and uncertain measurements. Each grid in this map is assigned a value that scores the probability of the grid space being occupied by obstacles. Semantic occupancy perception originates from SSCNet [24], which predicts the occupied status and semantics of all voxels in an indoor scene from a single image. However, studying occupancy perception in outdoor scenes is imperative for autonomous driving, as opposed to indoor scenes. MonoScene [25] is a pioneering work of outdoor scene occupancy perception using only a monocular camera. Contemporary with MonoScene, Tesla announced its brand-new camera-only occupancy network at the CVPR 2022 workshop on Autonomous Driving [26]. This new network comprehensively understands the 3D environment surrounding a vehicle according to surround-view RGB images. Subsequently, occupancy perception has attracted extensive attention, catalyzing a surge in research on occupancy perception for autonomous driving in recent years. The chronological overview in Fig. 2 indicates rapid development in occupancy perception since 2023.
@@ 2.3.1 Bird’s-Eye-View Perception
3D reconstruction is a traditional but important topic in the computer vision and robotics communities [62, 63, 64, 65]. The objective of 3D reconstruction from images is to construct 3D of an object or scene based on 2D images captured from one or more viewpoints. Early methods exploited shape-from-shading [66] or structure-from-motion [67]. Afterwards, the neural radiation field (NeRF) [68] introduced a novel paradigm for 3D reconstruction, which learned the density and color fields of 3D scenes, producing results with unprecedented detail and fidelity. However, such performance necessitates substantial training time and resources for rendering [69, 70, 71], especially for high-resolution output. Recently, 3D Gaussian splatting (3D GS) [72] has addressed this issue by redefining a paradigm-shifting approach to scene representation and rendering. Specifically, it represents scene representation with millions of 3D Gaussian functions in an explicit way, achieving faster and more efficient rendering [73]. 3D reconstruction emphasizes the geometric quality and visual appearance of the scene. In comparison, voxel-wise occupancy perception has lower resolution and visual appearance requirements, focusing instead on the occupancy distribution and semantic understanding of the scene.
@@ 3.1.1 General Pipeline
Figure 4: Architecture for LiDAR-centric occupancy perception: Solely the 2D branch [74, 78], solely the 3D branch [11, 27, 99], and integrating both 2D and 3D branches [29].
Fig. 4 illustrates the general pipeline of LiDAR-centric occupancy perception. The input point cloud first undergoes voxelization and feature extraction, followed by representation enhancement via an encoder-decoder module. Ultimately, the complete and dense occupancy of the scene is inferred. Specifically, given a point cloud $P\in\mathbb{R}^{N\times 3}$, we generate a series of initial voxels and extract their features. These voxels are distributed in a 3D volume [27, 80, 29, 99], a 2D BEV plane [74, 29], or three 2D tri-perspective view planes [78]. This operation constructs the 3D feature volume or 2D feature map, denoted as $V_{\text{init}-3D}\in\mathbb{R}^{X\times Y\times Z\times D}$ and $V_{\text{init}-2D}\in\mathbb{R}^{X\times Y\times D}$ respectively. $N$ represents the number of points; $\left(X,Y,Z\right)$ are the length, width, and height; $D$ mean the feature dimensions of voxels. In addition to voxelizing in regular Euclidean space, PointOcc [78] builds tri-perspective 2D feature maps in a cylindrical coordinate system. The cylindrical coordinate system aligns more closely with the spatial distribution of points in the LiDAR point cloud, where points closer to the LiDAR sensor are denser than those at farther distances. Therefore, it is reasonable to use smaller-sized cylindrical voxels for fine-grained modeling in nearby areas. The voxelization and feature extraction of point clouds can be formulated as:
$V_{\text{init}-2D\text{/}3D}=\Phi_{V}\left(\Psi_{V}\left(P\right)\right),$ (3)
where $\Psi_{V}$ stands for pillar or cubic voxelization. $\Phi_{V}$ is a feature extractor that extracts neural features of voxels (e.g., using PointPillars [100], VoxelNet [101], and MLP) [74, 78], or directly counts the geometric features of points within the voxel (e.g., mean, minimum, and maximum heights) [29, 99]. Encoder and decoder can be various modules to enhance features. The final 3D occupancy inference involves applying convolution [27, 29, 77] or MLP [74, 78, 80] on the enhanced features to infer the occupied status $\left\{1,0\right\}$ of each voxel, and even estimate its semantic category:
@@ 3.2.1 General Pipeline
Inspired by Tesla’s technology of the perception system for their autonomous vehicles [26], vision-centric occupancy perception has garnered significant attention both in industry and academia. Compared to LiDAR-centric methods, vision-centric occupancy perception, which solely relies on camera sensors, represents a current trend. There are three main reasons: (i) Cameras are cost-effective for large-scale deployment on vehicles. (ii) RGB images capture rich environmental textures, aiding in the understanding of scenes and objects such as traffic signs and lane lines. (iii) The burgeoning advancement of deep learning technologies enables a possibility to achieve 3D occupancy perception from 2D vision. Vision-centric occupancy perception can be divided into monocular solutions [102, 53, 25, 50, 51, 32, 54, 88, 84] and multi-camera solutions [52, 103, 30, 37, 60, 79, 104, 31, 81, 8]. Multi-camera perception, which covers a broader field of view, follows a general pipeline as shown in Fig. 5. It begins by extracting front-view feature maps from multi-camera images, followed by a 2D-to-3D transformation, spatial information fusion, and optional temporal information fusion, culminating with an occupancy head that infers the environmental 3D occupancy.
Figure 5: Architecture for vision-centric occupancy perception: Methods without temporal fusion [30, 37, 81, 75, 35, 31, 105, 87, 82, 83]; Methods with temporal fusion [79, 8, 55, 10, 85, 4].
@@ 3.2.3 Information Fusion in Vision-Centric Occupancy
In a multi-camera setting, each camera’s front-view feature map describes a part of the scene. To comprehensively understand the scene, it is necessary to spatially fuse the information from multiple feature maps. Additionally, objects in the scene might be occluded or in motion. Temporally fusing feature maps of multiple frames can help reason about the occluded areas and recognize the motion status of objects.
@@
Once the alignment is completed, the historical and current features are fed to the feature fusion module to enhance the representation, especially to strengthen the reasoning ability for occlusion and the recognition ability of moving objects. There are three main streamlines to feature fusion, namely convolution, cross attention, and adaptive mixing. PanoOcc [4] concatenates the previous features with the current ones, then fuses them using a set of 3D residual convolution blocks. Many occupancy perception methods [32, 55, 22, 84, 109] utilize cross attention for fusion. The process is similar to multi-camera cross attention (refer to Eq. 8), but the difference is that 3D-space voxels are projected to 2D multi-frame feature maps instead of multi-camera feature maps. Moreover, FullySparse [107] employs adaptive mixing [115] for temporal information fusion. For the query feature of the current frame, FullySparse samples $S_{n}$ features from historical frames, and aggregates them through adaptive mixing. Specifically, the sampled features are multiplied by the channel mixing matrix $W_{C}$ and the point mixing matrix $W_{S_{n}}$, respectively. These mixing matrices are dynamically generated from the query feature $F_{q}$ of the current frame:
@@
The features resulting from spatial and temporal information fusion are processed by various types of heads to determine 3D occupancy. These include convolutional heads, mask decoder heads, linear projection heads, and linear projection with threshold heads. Convolution-based heads [10, 60, 37, 103, 75, 25, 7] consist of multiple 3D convolutional layers. Mask decoder-based heads [85, 54, 107, 90], inspired by MaskFormer [116] and Mask2Former [117], formalize 3D semantic occupancy prediction into the estimation of a set of binary 3D masks, each associated with a corresponding semantic category. Specifically, they compute per-voxel embeddings and assess per-query embeddings along with their related semantics. The final occupancy predictions are obtained by calculating the dot product of these two embeddings. Linear projection-based heads [4, 31, 35, 50, 84, 32, 89] leverage lightweight MLPs on the dimension of feature channels to produce occupied status and semantics. Furthermore, for the occupancy methods[83, 81, 87, 105, 91] based on NeRF [68], their occupancy heads use two separate MLPs ($\text{MLP}_{\sigma}$, $\text{MLP}_{s}$) to estimate the density volume $V_{\sigma}$ and the semantic volume $V_{S}$. Then the occupied voxels are selected based on a given confidence threshold $\tau$, and their semantic categories are determined based on $V_{S}$:
@@
Figure 7: Architecture for multi-modal occupancy perception: Fusion of information from point clouds and images [11, 94, 12, 15, 95]. Dashed lines signify additional fusion of perspective-view feature maps [14]. $\odot$ represents element-wise product. $\delta$ is a learnable weight.